OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (-6, 4) and (2, 0). Answer y - 4 = -1/2 (x + 6) y - 4 = 2(x + 6) y + 6 = -1/2 (x - 4) y + 6 = 2(x - 4)

OpenStudy (anonymous):

we find slope Slope = Rise/Run = (0 - 4) / (2 + 6) = -4/8 = -1/2 using slope -1/2 and point (-6, 4) we get y - 4 = -1/2 (x + 6) as the correct choice

OpenStudy (callisto):

\[\frac{y-4}{x-(-6)} = \frac{0-4}{2-(-6)}\]\[\frac{y-4}{x+6} = \frac{-4}{8}\]\[{y-4} = -\frac{1}{2}{(x+6})\]